OpenStudy (anonymous):

completing the square x^2-3x-4=0

OpenStudy (anonymous):

im so stuck!

OpenStudy (anonymous):

\[x^{2}-3x-4=0

OpenStudy (anonymous):

( x+(1/6) )^2-(145/36)

OpenStudy (anonymous):

add 4 to both sides x^2 - 3x = 4 divide the coefficient of the base 1 by 2, then square it, and add to both sides. -1.5^2 = 2.25 x^2 - 3x + 2.25 = 4 + 2.25 (x - 1.5)^2 = 4 + 2.25 (x - 1.5)^2 = 6.25 take the square root of both sides x - 1.5 = +/- SqRt(6.25) ______________ x - 1.5 = 2.5 x = 4 x - 1.5 = -2.5 x = 1

OpenStudy (anonymous):

Thank you so much! Both of you big help!

OpenStudy (anonymous):

(4)^2 - 3(4) - 4 = 0 16 - 12 - 4 = 0 0 = 0 (-1)^2 - 3(-1) - 4 = 0 1 + 3 - 4 = 0 0 = 0 x = (4, -1) CORRECTION I WROTE 1, THE ANSWER IS REALLY -1 (AND 4)